OpenStudy (anonymous):

Prove that in each group, there is only one element of identity. That is, only one e such that e*a(where a is any element of the group) = a.

OpenStudy (anonymous):

Start out by assuming there is a second identity, i in the group. Show that i = e.

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

Assume that there is an f such that f*a = a. e*a = a is given Since f*a and e*a equal the same element, f*a = e*a [transitive law for equality]. Multiply both sides from the right by a^(-1) (the inverse of a): (r*a)*a^(-1) = (e*a)*a^(-1) f*(a*a*(-1)) = e*(a*a^(-1)) [associative property] f = e [inverse property] This proof hinges on the uniqueness of inverse for all elements in g.